Show steps to solve
0.105x+0.06(20,000)-x=1,740

Answer by srinivas.g(540)

0.105 x + 0.06 * 20,000 -x = 1740
0.105 x +1200 -x = 1740
-0.895 x +1200 = 1740
subtract 12 on both sides
-0.895 X +1200-1200 =1740-1200
-0.895 x = 540
divide with -0.895 on both sides
-0.895 x/-0.895 = 540/-0.895
x = -603.35